a square + b square = c square when b square = 5ft and c square = 10ft round to the nearest tenth
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

a^2+b^2=c^2
a^2+5^2=10^2
a^2+25=100
a^2=100-25
a^2=75
a=sqrt75
a=8.7 ans.

Answer by Alan3354(69443)

Do you mean b = 5 ft ?
And c = 10 ft ?
b^2 would be 5 square feet, not 5 ft